Solve the following equation for x: 3(x + 2) = 7x – 10
Can someone please help me with this equation. Thank you

Found 2 solutions by checkley77, scott8148:
Answer by checkley77(12844)

3(x + 2) = 7x – 10
3X+6=7X-10
3X-7X=-10-6
-4X=-16
X=-16/-4
X=4 ANSWER.
PROOF:
3(4+2)=7*4-10
3*6=28-10
18=28-10
18=18

Answer by scott8148(6628)

distributing the 3 __ 3x+6=7x-10

adding 10 __ 3x+16=7x

subtracting 3x __ 16=4x

dividing by 4 __ 4=x